### v2

@@ B.2.1 Without noise
B.2.2 With noise at level 5%
| Gini_generalized | 1.1 | 0.5391 | 0.4359 | 4.92 | 0.6306 | 0.6334 | 2.40 | 0.5501 | 0.5470 | 2.32 | 0.5959 | 0.3566 | 2.44 | 0.6283 | 0.4736 | 1.82 | 0.5561 | 0.3499 | 2.30 | 0.2688 | 0.1364 | 1.58 | 0.7066 | 0.0552 |
| Gini | - | 0.3996 | 0.3447 | - | 0.5779 | 0.5773 | - | 0.5140 | 0.5140 | - | 0.5373 | 0.3633 | - | 0.5660 | 0.4336 | - | 0.5344 | 0.3432 | - | 0.1457 | 0.1341 | - | 0.6802 | 0.1054 |
| Euclidean | - | 0.3344 | 0.2906 | - | 0.6070 | 0.6115 | - | 0.5120 | 0.5120 | - | 0.5136 | 0.2574 | - | 0.5975 | 0.4576 | - | 0.5479 | 0.3335 | - | 0.5428 | 0.1250 | - | 0.7150 | 0.0155 |
| Manhattan | - | 0.3792 | 0.3276 | - | 0.6071 | 0.6115 | - | 0.5110 | 0.5110 | - | 0.4994 | 0.2540 | - | 0.6476 | 0.4912 | - | 0.5332 | 0.3335 | - | 0.5508 | 0.0977 | - | 0.7150 | 0.0155 |
| Minkowski | 3.0 | 0.3309 | 0.2878 | 3.0 | 0.6072 | 0.6115 | 3.0 | 0.5120 | 0.5120 | 3.0 | 0.5045 | 0.2607 | 3.0 | 0.6127 | 0.4752 | 3.0 | 0.5515 | 0.3204 | 3.0 | 0.5340 | 0.1409 | 3.0 | 0.7150 | 0.0155 |
| Cosine | - | 0.4291 | 0.3645 | - | 0.5976 | 0.6006 | - | 0.4885 | 0.4900 | - | 0.5169 | 0.2708 | - | 0.6354 | 0.4880 | - | 0.5452 | 0.3627 | - | 0.5901 | 0.4909 | - | 0.6305 | 0.1640 |
| Canberra | - | 0.5411 | 0.4386 | - | 0.5980 | 0.6021 | - | 0.5439 | 0.5300 | - | 0.4764 | 0.1987 | - | 0.6047 | 0.4592 | - | 0.5521 | 0.3920 | - | 0.2220 | 0.1068 | - | 0.6860 | 0.1745 |
| Hellinger | - | 0.7727 | 0.6640 | - | 0.5949 | 0.5860 | - | 0.4916 | 0.4900 | - | 0.4628 | 0.2242 | - | 0.4992 | 0.3440 | - | 0.5426 | 0.4308 | - | 0.2941 | 0.2136 | - | 0.7150 | 0.0432 |
| Jensen-Shannon | - | 0.1289 | 0.3590 | - | 0.3085 | 0.5554 | - | 0.2500 | 0.5000 | - | 0.1495 | 0.0132 | - | 0.5386 | 0.1872 | - | 0.5328 | 0.0918 | - | 0.2626 | 0.1409 | - | 0.7150 | 0.0518 |
| PearsonChi2 | - | 0.5953 | 0.5153 | - | 0.5111 | 0.5109 | - | 0.5030 | 0.5030 | - | 0.4480 | 0.2309 | - | 0.4977 | 0.4224 | - | 0.4466 | 0.2981 | - | 0.5010 | 0.4909 | - | 0.6051 | 0.2381 |
| VicisSymmetric1 | - | 0.5344 | 0.4727 | - | 0.5036 | 0.5014 | - | 0.4917 | 0.5070 | - | 0.4717 | 0.2345 | - | 0.4313 | 0.3744 | - | 0.5110 | 0.3162 | - | 0.3415 | 0.2864 | - | 0.5583 | 0.2870 |
| Hassanat | - | 0.4426 | 0.3760 | - | 0.5572 | 0.5539 | - | 0.5619 | 0.5610 | - | 0.5826 | 0.3333 | - | 0.5862 | 0.4496 | - | 0.5901 | 0.5520 | - | 0.3647 | 0.3091 | - | 0.6915 | 0.1486 |
| Gini_generalized | 2.68 | 0.0939 | 0.1067 | 4.56 | 0.2085 | 0.2021 | 2.24 | 0.9283 | 0.9245 | 2.04 | 0.5750 | 0.5722 | 2.30 | 0.3871 | 0.2850 | 4.76 | 0.3050 | 0.2872 | 1.40 | 0.7677 | 0.6000 | 1.50 | 0.0918 | 0.0514 |
| Gini | - | 0.0761 | 0.1000 | - | 0.1207 | 0.1010 | - | 0.9107 | 0.9034 | - | 0.5460 | 0.5431 | - | 0.3542 | 0.2684 | - | 0.2654 | 0.2576 | - | 0.7086 | 0.6145 | - | 0.0774 | 0.0421 |
| Euclidean | - | 0.0728 | 0.0933 | - | 0.0610 | 0.0449 | - | 0.8771 | 0.8489 | - | 0.5448 | 0.5382 | - | 0.3563 | 0.2337 | - | 0.2105 | 0.2292 | - | 0.4877 | 0.5594 | - | 0.0358 | 0.0188 |
| Manhattan | - | 0.0761 | 0.1000 | - | 0.0554 | 0.0394 | - | 0.8735 | 0.8437 | - | 0.5621 | 0.5575 | - | 0.3669 | 0.2632 | - | 0.2055 | 0.2221 | - | 0.4877 | 0.5594 | - | 0.0455 | 0.0235 |
| Minkowski | 3.0 | 0.0690 | 0.0867 | 3.0 | 0.0610 | 0.0449 | 3.0 | 0.8793 | 0.8524 | 3.0 | 0.5331 | 0.5240 | 3.0 | 0.3552 | 0.2325 | 3.0 | 0.2098 | 0.2292 | 3.0 | 0.4877 | 0.5594 | 3.0 | 0.0336 | 0.0188 |
| Cosine | - | 0.0211 | 0.0200 | - | 0.1536 | 0.1235 | - | 0.9131 | 0.9034 | - | 0.5460 | 0.5431 | - | 0.3990 | 0.3081 | - | 0.1968 | 0.2162 | - | 0.6776 | 0.6725 | - | 0.0554 | 0.0467 |
| Canberra | - | 0.0539 | 0.0667 | - | 0.3879 | 0.3483 | - | 0.9223 | 0.9209 | - | 0.6073 | 0.6062 | - | 0.4330 | 0.3402 | - | 0.2386 | 0.2304 | - | 0.5703 | 0.5913 | - | 0.0806 | 0.0470 |
| Hellinger | - | 0.1529 | 0.1533 | - | 0.4685 | 0.4213 | - | 0.7039 | 0.6893 | - | 0.6826 | 0.5002 | - | 0.4312 | 0.3348 | - | 0.4326 | 0.3630 | - | 0.6697 | 0.5710 | - | 0.2237 | 0.1401 |
| Jensen-Shannon | - | 0.4233 | 0.4467 | - | 0.0490 | 0.1183 | - | 0.1380 | 0.3691 | - | 0.2849 | 0.5337 | - | 0.3487 | 0.5905 | - | 0.3686 | 0.4350 | - | 0.3864 | 0.5623 | - | 0.0000 | 0.0000 |
| PearsonChi2 | - | 0.2017 | 0.1867 | - | 0.3128 | 0.3830 | - | 0.8853 | 0.8647 | - | 0.4072 | 0.4086 | - | 0.4221 | 0.3261 | - | 0.2652 | 0.2351 | - | 0.7337 | 0.5855 | - | 0.2041 | 0.1728 |
| VicisSymmetric1 | - | 0.0667 | 0.0267 | - | 0.4796 | 0.4622 | - | 0.5351 | 0.5009 | - | 0.5330 | 0.5283 | - | 0.4063 | 0.2567 | - | 0.3746 | 0.3275 | - | 0.3966 | 0.4986 | - | 0.2229 | 0.1353 |
| Hassanat | - | 0.0637 | 0.0800 | - | 0.3990 | 0.3541 | - | 0.9245 | 0.9157 | - | 0.5578 | 0.5527 | - | 0.3962 | 0.2581 | - | 0.3502 | 0.2943 | - | 0.7850 | 0.7710 | - | 0.0351 | 0.0234 |
@@ B.2.2 With noise at level 5%
B.2.3 With noise at level 10%
| Gini_generalized | 3.38 | 0.0946 | 0.1067 | 4.78 | 0.2116 | 0.2022 | 2.24 | 0.9283 | 0.9245 | 2.94 | 0.5864 | 0.5815 | 2.14 | 0.3871 | 0.2863 | 4.44 | 0.3080 | 0.2872 | 1.34 | 0.7672 | 0.5986 | 2.38 | 0.1010 | 0.0560 |
| Gini | - | 0.0768 | 0.1000 | - | 0.1207 | 0.1010 | - | 0.9107 | 0.9034 | - | 0.5586 | 0.5527 | - | 0.3614 | 0.2722 | - | 0.2651 | 0.2576 | - | 0.7101 | 0.6159 | - | 0.0356 | 0.0186 |
| Euclidean | - | 0.0806 | 0.1000 | - | 0.0610 | 0.0449 | - | 0.8771 | 0.8489 | - | 0.5536 | 0.5479 | - | 0.3571 | 0.2350 | - | 0.2097 | 0.2281 | - | 0.4877 | 0.5594 | - | 0.0435 | 0.0235 |
| Manhattan | - | 0.0761 | 0.1000 | - | 0.0554 | 0.0394 | - | 0.8735 | 0.8437 | - | 0.5488 | 0.5432 | - | 0.3640 | 0.2581 | - | 0.2055 | 0.2221 | - | 0.4877 | 0.5594 | - | 0.0502 | 0.0281 |
| Minkowski | 3.0 | 0.0751 | 0.0933 | 3.0 | 0.0610 | 0.0449 | 3.0 | 0.8793 | 0.8524 | 3.0 | 0.5742 | 0.5672 | 3.0 | 0.3551 | 0.2325 | 3.0 | 0.2105 | 0.2304 | 3.0 | 0.4877 | 0.5594 | 3.0 | 0.0435 | 0.0235 |
| Cosine | - | 0.0216 | 0.0200 | - | 0.1514 | 0.1235 | - | 0.9131 | 0.9034 | - | 0.5523 | 0.5479 | - | 0.3989 | 0.3056 | - | 0.1968 | 0.2162 | - | 0.6776 | 0.6725 | - | 0.0626 | 0.0467 |
| Canberra | - | 0.0303 | 0.0333 | - | 0.3879 | 0.3484 | - | 0.9223 | 0.9209 | - | 0.5474 | 0.5483 | - | 0.4397 | 0.3543 | - | 0.2333 | 0.2257 | - | 0.5599 | 0.5696 | - | 0.1254 | 0.0699 |
| Hellinger | - | 0.3318 | 0.3600 | - | 0.5159 | 0.4602 | - | 0.7039 | 0.6893 | - | 0.5904 | 0.4710 | - | 0.4265 | 0.3416 | - | 0.4041 | 0.3617 | - | 0.6697 | 0.5710 | - | 0.2508 | 0.1355 |
| Jensen-Shannon | - | 0.4514 | 0.4800 | - | 0.0525 | 0.1295 | - | 0.1388 | 0.3726 | - | 0.2849 | 0.5337 | - | 0.4401 | 0.5661 | - | 0.3701 | 0.4373 | - | 0.4003 | 0.5667 | - | 0.0000 | 0.0000 |
| PearsonChi2 | - | 0.2908 | 0.2667 | - | 0.2757 | 0.2408 | - | 0.8904 | 0.8683 | - | 0.4979 | 0.4906 | - | 0.5424 | 0.4004 | - | 0.3915 | 0.2918 | - | 0.7548 | 0.5826 | - | 0.2871 | 0.1543 |
| VicisSymmetric1 | - | 0.0000 | 0.0000 | - | 0.2393 | 0.2359 | - | 0.5108 | 0.4905 | - | 0.4841 | 0.4764 | - | 0.4278 | 0.3028 | - | 0.3789 | 0.3238 | - | 0.3890 | 0.4841 | - | 0.2037 | 0.1351 |
| Hassanat | - | 0.0457 | 0.0533 | - | 0.3931 | 0.3484 | - | 0.9237 | 0.9174 | - | 0.5488 | 0.5432 | - | 0.3917 | 0.2568 | - | 0.3494 | 0.2919 | - | 0.7838 | 0.7696 | - | 0.0629 | 0.0374 |
| Gini_generalized | 1.1 | 0.4815 | 0.4045 | 4.98 | 0.6305 | 0.6334 | 2.38 | 0.5510 | 0.5480 | 2.32 | 0.5966 | 0.3532 | 2.16 | 0.6248 | 0.4608 | 1.74 | 0.5615 | 0.3435 | 2.30 | 0.2688 | 0.1364 | 1.56 | 0.7066 | 0.0552 |
| Gini | - | 0.3924 | 0.3389 | - | 0.5787 | 0.5780 | - | 0.5140 | 0.5140 | - | 0.5389 | 0.3601 | - | 0.5608 | 0.4336 | - | 0.5451 | 0.3433 | - | 0.1457 | 0.1341 | - | 0.6802 | 0.1054 |
| Euclidean | - | 0.3351 | 0.2905 | - | 0.6032 | 0.6079 | - | 0.5120 | 0.5120 | - | 0.5136 | 0.2574 | - | 0.6109 | 0.4688 | - | 0.5425 | 0.3302 | - | 0.5428 | 0.1250 | - | 0.7150 | 0.0155 |
| Manhattan | - | 0.3825 | 0.3304 | - | 0.6081 | 0.6123 | - | 0.5120 | 0.5120 | - | 0.5040 | 0.2540 | - | 0.6367 | 0.4768 | - | 0.5332 | 0.3335 | - | 0.5508 | 0.0977 | - | 0.7150 | 0.0155 |
| Minkowski | 3.0 | 0.3342 | 0.2906 | 3.0 | 0.6026 | 0.6072 | 3.0 | 0.5120 | 0.5120 | 3.0 | 0.5018 | 0.2607 | 3.0 | 0.6067 | 0.4864 | 3.0 | 0.5529 | 0.3237 | 3.0 | 0.5340 | 0.1409 | 3.0 | 0.7150 | 0.0155 |
| Cosine | - | 0.4329 | 0.3673 | - | 0.5976 | 0.6006 | - | 0.4885 | 0.4900 | - | 0.5169 | 0.2708 | - | 0.6290 | 0.4800 | - | 0.5452 | 0.3627 | - | 0.5901 | 0.4909 | - | 0.6305 | 0.1640 |
| Canberra | - | 0.4749 | 0.3959 | - | 0.5967 | 0.6013 | - | 0.6016 | 0.5900 | - | 0.6551 | 0.2808 | - | 0.5712 | 0.4352 | - | 0.5444 | 0.3854 | - | 0.2182 | 0.1023 | - | 0.6871 | 0.1797 |
| Hellinger | - | 0.6010 | 0.5212 | - | 0.5944 | 0.5860 | - | 0.4828 | 0.4840 | - | 0.6006 | 0.1911 | - | 0.4499 | 0.3632 | - | 0.5125 | 0.3472 | - | 0.3535 | 0.2909 | - | 0.7150 | 0.0432 |
| Jensen-Shannon | - | 0.1289 | 0.3590 | - | 0.3085 | 0.5554 | - | 0.2500 | 0.5000 | - | 0.0001 | 0.0066 | - | 0.6888 | 0.2576 | - | 0.4546 | 0.0754 | - | 0.2999 | 0.1477 | - | 0.7150 | 0.0535 |
| PearsonChi2 | - | 0.5366 | 0.4931 | - | 0.4841 | 0.4753 | - | 0.4909 | 0.4910 | - | 0.5051 | 0.2276 | - | 0.5122 | 0.3776 | - | 0.5686 | 0.4965 | - | 0.4801 | 0.4818 | - | 0.5753 | 0.3588 |
| VicisSymmetric1 | - | 0.4569 | 0.4017 | - | 0.5169 | 0.5131 | - | 0.4917 | 0.5070 | - | 0.4001 | 0.2474 | - | 0.4995 | 0.4064 | - | 0.5043 | 0.3392 | - | 0.3655 | 0.3318 | - | 0.6194 | 0.2139 |
| Hassanat | - | 0.4245 | 0.3618 | - | 0.5560 | 0.5525 | - | 0.5352 | 0.5350 | - | 0.6532 | 0.3668 | - | 0.6169 | 0.4784 | - | 0.5901 | 0.5520 | - | 0.3596 | 0.3000 | - | 0.6912 | 0.1520 |

### v3

@@ 6 Conclusion
With noise at level 5%
Ionosphere Banknote German Heart Balance Haberman Wholesale Indian Liver Patient Distance #P #R #P #R #P #R #P #R #P #R #P #R #P #R #P #R Gini_generalized 1.1 0.5391 0.4359 4.92 0.6306 0.6334 2.40 0.5501 0.5470 2.32 0.5959 0.3566 2.44 0.6283 0.4736 1.82 0.5561 0.3499 2.30 0.2688 0.1364 1.58 0.7066 0.0552 Gini - 0.3996 0.3447 - 0.5779 0.5773 - 0.5140 0.5140 - 0.5373 0.3633 - 0.5660 0.4336 - 0.5344 0.3432 - 0.1457 0.1341 - 0.6802 0.1054 Euclidean - 0.3344 0.2906 - 0.6070 0.6115 - 0.5120 0.5120 - 0.5136 0.2574 - 0.5975 0.4576 - 0.5479 0.3335 - 0.5428 0.1250 - 0.7150 0.0155 Manhattan - 0.3792 0.3276 - 0.6071 0.6115 - 0.5110 0.5110 - 0.4994 0.2540 - 0.6476 0.4912 - 0.5332 0.3335 - 0.5508 0.0977 - 0.7150 0.0155 Minkowski 3.0 0.3309 0.2878 3.0 0.6072 0.6115 3.0 0.5120 0.5120 3.0 0.5045 0.2607 3.0 0.6127 0.4752 3.0 0.5515 0.3204 3.0 0.5340 0.1409 3.0 0.7150 0.0155 Cosine - 0.4291 0.3645 - 0.5976 0.6006 - 0.4885 0.4900 - 0.5169 0.2708 - 0.6354 0.4880 - 0.5452 0.3627 - 0.5901 0.4909 - 0.6305 0.1640 Canberra - 0.5411 0.4386 - 0.5980 0.6021 - 0.5439 0.5300 - 0.4764 0.1987 - 0.6047 0.4592 - 0.5521 0.3920 - 0.2220 0.1068 - 0.6860 0.1745 Hellinger - 0.7727 0.6640 - 0.5949 0.5860 - 0.4916 0.4900 - 0.4628 0.2242 - 0.4992 0.3440 - 0.5426 0.4308 - 0.2941 0.2136 - 0.7150 0.0432 Jensen-Shannon - 0.1289 0.3590 - 0.3085 0.5554 - 0.2500 0.5000 - 0.1495 0.0132 - 0.5386 0.1872 - 0.5328 0.0918 - 0.2626 0.1409 - 0.7150 0.0518 PearsonChi2 - 0.5953 0.5153 - 0.5111 0.5109 - 0.5030 0.5030 - 0.4480 0.2309 - 0.4977 0.4224 - 0.4466 0.2981 - 0.5010 0.4909 - 0.6051 0.2381 VicisSymmetric1 - 0.5344 0.4727 - 0.5036 0.5014 - 0.4917 0.5070 - 0.4717 0.2345 - 0.4313 0.3744 - 0.5110 0.3162 - 0.3415 0.2864 - 0.5583 0.2870 Hassanat - 0.4426 0.3760 - 0.5572 0.5539 - 0.5619 0.5610 - 0.5826 0.3333 - 0.5862 0.4496 - 0.5901 0.5520 - 0.3647 0.3091 - 0.6915 0.1486
Iris Wine Breast-Cancer Sonar QSAR Vehicle Australian Glass Distance #P #R #P #R #P #R #P #R #P #R #P #R #P #R #P #R Gini_generalized 2.68 0.0939 0.1067 4.56 0.2085 0.2021 2.24 0.9283 0.9245 2.04 0.5750 0.5722 2.30 0.3871 0.2850 4.76 0.3050 0.2872 1.40 0.7677 0.6000 1.50 0.0918 0.0514 Gini - 0.0761 0.1000 - 0.1207 0.1010 - 0.9107 0.9034 - 0.5460 0.5431 - 0.3542 0.2684 - 0.2654 0.2576 - 0.7086 0.6145 - 0.0774 0.0421 Euclidean - 0.0728 0.0933 - 0.0610 0.0449 - 0.8771 0.8489 - 0.5448 0.5382 - 0.3563 0.2337 - 0.2105 0.2292 - 0.4877 0.5594 - 0.0358 0.0188 Manhattan - 0.0761 0.1000 - 0.0554 0.0394 - 0.8735 0.8437 - 0.5621 0.5575 - 0.3669 0.2632 - 0.2055 0.2221 - 0.4877 0.5594 - 0.0455 0.0235 Minkowski 3.0 0.0690 0.0867 3.0 0.0610 0.0449 3.0 0.8793 0.8524 3.0 0.5331 0.5240 3.0 0.3552 0.2325 3.0 0.2098 0.2292 3.0 0.4877 0.5594 3.0 0.0336 0.0188 Cosine - 0.0211 0.0200 - 0.1536 0.1235 - 0.9131 0.9034 - 0.5460 0.5431 - 0.3990 0.3081 - 0.1968 0.2162 - 0.6776 0.6725 - 0.0554 0.0467 Canberra - 0.0539 0.0667 - 0.3879 0.3483 - 0.9223 0.9209 - 0.6073 0.6062 - 0.4330 0.3402 - 0.2386 0.2304 - 0.5703 0.5913 - 0.0806 0.0470 Hellinger - 0.1529 0.1533 - 0.4685 0.4213 - 0.7039 0.6893 - 0.6826 0.5002 - 0.4312 0.3348 - 0.4326 0.3630 - 0.6697 0.5710 - 0.2237 0.1401 Jensen-Shannon - 0.4233 0.4467 - 0.0490 0.1183 - 0.1380 0.3691 - 0.2849 0.5337 - 0.3487 0.5905 - 0.3686 0.4350 - 0.3864 0.5623 - 0.0000 0.0000 PearsonChi2 - 0.2017 0.1867 - 0.3128 0.3830 - 0.8853 0.8647 - 0.4072 0.4086 - 0.4221 0.3261 - 0.2652 0.2351 - 0.7337 0.5855 - 0.2041 0.1728 VicisSymmetric1 - 0.0667 0.0267 - 0.4796 0.4622 - 0.5351 0.5009 - 0.5330 0.5283 - 0.4063 0.2567 - 0.3746 0.3275 - 0.3966 0.4986 - 0.2229 0.1353 Hassanat - 0.0637 0.0800 - 0.3990 0.3541 - 0.9245 0.9157 - 0.5578 0.5527 - 0.3962 0.2581 - 0.3502 0.2943 - 0.7850 0.7710 - 0.0351 0.0234
@@
With noise at level 10%
Iris Wine Breast-Cancer Sonar QSAR Vehicle Australian Glass Distance #P #R #P #R #P #R #P #R #P #R #P #R #P #R #P #R Gini_generalized 3.38 0.0946 0.1067 4.78 0.2116 0.2022 2.24 0.9283 0.9245 2.94 0.5864 0.5815 2.14 0.3871 0.2863 4.44 0.3080 0.2872 1.34 0.7672 0.5986 2.38 0.1010 0.0560 Gini - 0.0768 0.1000 - 0.1207 0.1010 - 0.9107 0.9034 - 0.5586 0.5527 - 0.3614 0.2722 - 0.2651 0.2576 - 0.7101 0.6159 - 0.0356 0.0186 Euclidean - 0.0806 0.1000 - 0.0610 0.0449 - 0.8771 0.8489 - 0.5536 0.5479 - 0.3571 0.2350 - 0.2097 0.2281 - 0.4877 0.5594 - 0.0435 0.0235 Manhattan - 0.0761 0.1000 - 0.0554 0.0394 - 0.8735 0.8437 - 0.5488 0.5432 - 0.3640 0.2581 - 0.2055 0.2221 - 0.4877 0.5594 - 0.0502 0.0281 Minkowski 3.0 0.0751 0.0933 3.0 0.0610 0.0449 3.0 0.8793 0.8524 3.0 0.5742 0.5672 3.0 0.3551 0.2325 3.0 0.2105 0.2304 3.0 0.4877 0.5594 3.0 0.0435 0.0235 Cosine - 0.0216 0.0200 - 0.1514 0.1235 - 0.9131 0.9034 - 0.5523 0.5479 - 0.3989 0.3056 - 0.1968 0.2162 - 0.6776 0.6725 - 0.0626 0.0467 Canberra - 0.0303 0.0333 - 0.3879 0.3484 - 0.9223 0.9209 - 0.5474 0.5483 - 0.4397 0.3543 - 0.2333 0.2257 - 0.5599 0.5696 - 0.1254 0.0699 Hellinger - 0.3318 0.3600 - 0.5159 0.4602 - 0.7039 0.6893 - 0.5904 0.4710 - 0.4265 0.3416 - 0.4041 0.3617 - 0.6697 0.5710 - 0.2508 0.1355 Jensen-Shannon - 0.4514 0.4800 - 0.0525 0.1295 - 0.1388 0.3726 - 0.2849 0.5337 - 0.4401 0.5661 - 0.3701 0.4373 - 0.4003 0.5667 - 0.0000 0.0000 PearsonChi2 - 0.2908 0.2667 - 0.2757 0.2408 - 0.8904 0.8683 - 0.4979 0.4906 - 0.5424 0.4004 - 0.3915 0.2918 - 0.7548 0.5826 - 0.2871 0.1543 VicisSymmetric1 - 0.0000 0.0000 - 0.2393 0.2359 - 0.5108 0.4905 - 0.4841 0.4764 - 0.4278 0.3028 - 0.3789 0.3238 - 0.3890 0.4841 - 0.2037 0.1351 Hassanat - 0.0457 0.0533 - 0.3931 0.3484 - 0.9237 0.9174 - 0.5488 0.5432 - 0.3917 0.2568 - 0.3494 0.2919 - 0.7838 0.7696 - 0.0629 0.0374
Ionosphere Banknote German Heart Balance Haberman Wholesale Indian Liver Patient Distance #P #R #P #R #P #R #P #R #P #R #P #R #P #R #P #R Gini_generalized 1.1 0.4815 0.4045 4.98 0.6305 0.6334 2.38 0.5510 0.5480 2.32 0.5966 0.3532 2.16 0.6248 0.4608 1.74 0.5615 0.3435 2.30 0.2688 0.1364 1.56 0.7066 0.0552 Gini - 0.3924 0.3389 - 0.5787 0.5780 - 0.5140 0.5140 - 0.5389 0.3601 - 0.5608 0.4336 - 0.5451 0.3433 - 0.1457 0.1341 - 0.6802 0.1054 Euclidean - 0.3351 0.2905 - 0.6032 0.6079 - 0.5120 0.5120 - 0.5136 0.2574 - 0.6109 0.4688 - 0.5425 0.3302 - 0.5428 0.1250 - 0.7150 0.0155 Manhattan - 0.3825 0.3304 - 0.6081 0.6123 - 0.5120 0.5120 - 0.5040 0.2540 - 0.6367 0.4768 - 0.5332 0.3335 - 0.5508 0.0977 - 0.7150 0.0155 Minkowski 3.0 0.3342 0.2906 3.0 0.6026 0.6072 3.0 0.5120 0.5120 3.0 0.5018 0.2607 3.0 0.6067 0.4864 3.0 0.5529 0.3237 3.0 0.5340 0.1409 3.0 0.7150 0.0155 Cosine - 0.4329 0.3673 - 0.5976 0.6006 - 0.4885 0.4900 - 0.5169 0.2708 - 0.6290 0.4800 - 0.5452 0.3627 - 0.5901 0.4909 - 0.6305 0.1640 Canberra - 0.4749 0.3959 - 0.5967 0.6013 - 0.6016 0.5900 - 0.6551 0.2808 - 0.5712 0.4352 - 0.5444 0.3854 - 0.2182 0.1023 - 0.6871 0.1797 Hellinger - 0.6010 0.5212 - 0.5944 0.5860 - 0.4828 0.4840 - 0.6006 0.1911 - 0.4499 0.3632 - 0.5125 0.3472 - 0.3535 0.2909 - 0.7150 0.0432 Jensen-Shannon - 0.1289 0.3590 - 0.3085 0.5554 - 0.2500 0.5000 - 0.0001 0.0066 - 0.6888 0.2576 - 0.4546 0.0754 - 0.2999 0.1477 - 0.7150 0.0535 PearsonChi2 - 0.5366 0.4931 - 0.4841 0.4753 - 0.4909 0.4910 - 0.5051 0.2276 - 0.5122 0.3776 - 0.5686 0.4965 - 0.4801 0.4818 - 0.5753 0.3588 VicisSymmetric1 - 0.4569 0.4017 - 0.5169 0.5131 - 0.4917 0.5070 - 0.4001 0.2474 - 0.4995 0.4064 - 0.5043 0.3392 - 0.3655 0.3318 - 0.6194 0.2139 Hassanat - 0.4245 0.3618 - 0.5560 0.5525 - 0.5352 0.5350 - 0.6532 0.3668 - 0.6169 0.4784 - 0.5901 0.5520 - 0.3596 0.3000 - 0.6912 0.1520